how much solution of 20% acid, must be added to a solution of 35% acid to get 60 gallons of a 25% solution?
Answer by lwsshak3(11628)

how much solution of 20% acid, must be added to a solution of 35% acid to get 60 gallons of a 25% solution?
***
let x=amt of 20% solution must be added
60-x=amt of 35% solution must be added
amt of acid on left side=amt of acid on right side
..
20%x+35%(60-x)=25%*60
.20x+21-.35x=15
.15x=6
x=40
60-x=20
amt of 20% solution must be added=40 gal.
amt of 35% solution must be added=20 gal.